OpenStudy (anonymous):

ln(5+4x)-ln(3+x)=ln3

myininaya (myininaya):

ln(x)-ln(y)=ln(x/y)

myininaya (myininaya):

Write ln[(5+4x)/{3+x)]=ln3

myininaya (myininaya):

base e both sides to get rid of the natural logs

myininaya (myininaya):

[(5+4x)/(3+x)]=3

myininaya (myininaya):

Solve for x

myininaya (myininaya):

5+4x=3(3+x) I multiply 3+x on both sides

myininaya (myininaya):

5+4x=9+3x Distributive property

myininaya (myininaya):

5+x=9 so x=9-5=4

myininaya (myininaya):

Always check your solution when working with log. You never know the solution may not be in the domain

OpenStudy (anonymous):

thank you for your help

myininaya (myininaya):

In this case it is, because 5+4x and 3+x is positive for x=4

myininaya (myininaya):

no problem

OpenStudy (anonymous):

I have one more question. Is that ok?

myininaya (myininaya):

Go ahead

OpenStudy (anonymous):

r=p-k ln t, for t

myininaya (myininaya):

So we want to solve for t. First get the ln(t) by itself

myininaya (myininaya):

Subtract p on both sides: r-p=-klnt

myininaya (myininaya):

Now divide -k on both sides: (r-p)/(-k)=lnt

myininaya (myininaya):

Base e both sides to get ride of the natural log

myininaya (myininaya):

e^[-(r-p)/k]=t